OpenStudy (anonymous):

what is the LCM of 220 and 400?

OpenStudy (anonymous):

4400

OpenStudy (anonymous):

LCM = Least Common Multiple = smallest number divisible cleanly by both numbers Prime factorization of 220 = 2*2*5*11 Prime factorization of 400 = 2*2*2*2*5*5 Therefore, the LCM is 2*2*2*2*5*5*11 = 4400

OpenStudy (anonymous):

thanks